OpenStudy (anonymous):

topic: integral calculus, help me to integrate this pleasee, Integral of (2 + 5/x)^2 dx

OpenStudy (anonymous):

\[\int\limits (2 + 5/x)^2 dx\]

OpenStudy (anonymous):

@iGreen @MARC_ @nikvist

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ilo

OpenStudy (anonymous):

@Ilovecake

OpenStudy (irishboy123):

simplest is just expand the brackets and do each term

OpenStudy (anonymous):

so it will become (4+ 25/x^2 ?

OpenStudy (anonymous):

@Ilovecake

OpenStudy (ilovecake):

Yes

OpenStudy (anonymous):

then

OpenStudy (irishboy123):

no no no

OpenStudy (anonymous):

how? @IrishBoy123

OpenStudy (irishboy123):

(2+ 5/x) times (2+ 5/x)

OpenStudy (anonymous):

ah , what should i do next

OpenStudy (irishboy123):

do that expansion then integrate each bit separately.

OpenStudy (anonymous):

i will multiply it each by piece?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

yep post yr answer here

OpenStudy (anonymous):

( 4 + 10/x+10/x +4)

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (phi):

is the original expression \[ \left(2 + \frac{5}{x}\right)^2 \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

after squaring did you get \[ 4 + \frac{25}{x^2} + \frac{20}{x} \]

OpenStudy (anonymous):

yes then what will i do

OpenStudy (phi):

write the problem as \[ 4\int \ dx + 25\int x^{-2} \ dx + 20 \int \frac{dx}{x} \]

OpenStudy (phi):

I assume you can do some of those integrals ?

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (phi):

are you saying you don't know how to do \[ 4 \int dx \] ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

the integral of dx is x based on this question, they expect you to know quite a bit of integral calculus, and integral of dx is the first baby step, which means you have to learn a lot of stuff to do this problem. And it's very hard to learn or teach calculus typing into this box. Maybe watch Khan's videos https://www.khanacademy.org/math/integral-calculus/indefinite-definite-integrals/indefinite_integrals/v/antiderivatives-and-indefinite-integrals